OpenStudy (anonymous):

ABCD is a rectangle if ac = 3c/9 and BD = 4- c, find the length of each diagonal a.) 2, 2 b.) 1, 1 c.) 3, 3 d.) 16, 16

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

ok i got it ready

OpenStudy (anonymous):

okay , i am ready :3

OpenStudy (anonymous):

ABCD is a rectangle if ac = 3c/9 and BD = 4- c, find the length of each diagonal a.) 2, 2 b.) 1, 1 c.) 3, 3 d.) 16, 16 ok first plug in the 1 where the c is you multiply the 3*1/9 and get 3/9 then you plug it in on the other side 4-1= 3 they are not the same so continue to plug them in until u det one with the same number and i will wait to check ur answer

OpenStudy (anonymous):

now try 2 instead of 1

OpenStudy (anonymous):

3*2/9 and you get 2/9 4-2=2 ^^ like this?

OpenStudy (anonymous):

6/9 *

OpenStudy (anonymous):

3*3/9= 9/9 4-3=1 ?

OpenStudy (anonymous):

great they are not matching so lets try 3

OpenStudy (anonymous):

wat does 9/9 reduce to

OpenStudy (anonymous):

1 = 1 .. so they are the same :D

OpenStudy (anonymous):

and the answer would be B ?

OpenStudy (anonymous):

great i give u gold star :) if you are telling the diag yes if u are looking for c ith would be c but you are correct it is b

OpenStudy (anonymous):

u get?:)

OpenStudy (saranya):

Use the Pythagorean Theorem a² + b² = c² to determine the diagonal, treating c as your diagonal..

OpenStudy (anonymous):

i do thank you guys :)